# ELMM: Efficient Lightweight Multimodal Large Language Models for Multimodal Knowledge Graph Completion

### v1

@@ 4.5 Parameter Sensitivity Analysis
Figure 4: Hyperparameter analysis results. Left: Relationship between Hits@1 and the number of attention pruning layers. Right: Relationship between Hits@1 and learning rate.
@@ 4.5.1 Number of Pruning Layers
4.5.2 Learning Rate
[…]
@@ 4.6 Impact of Backbone MLLMs on MKGC Performance
To investigate whether ELMM can continuously benefit from the rapid evolution of multimodal large language models (MLLMs), we study the impact of different backbone models on completion performance. Specifically, we select LLaVA-1.5 7B Liu et al. (2024), LLaVA-1.5 13B Liu et al. (2024), and Qwen2.5-VL 7B Bai et al. (2025a) as comparative models. The results are presented in Table 4.
@@
LLaVA-1.5 7B 112 34.1 47.3 60.7
LLaVA-1.5 13B 108 34.8 47.7 61.0
Qwen2.5-VL 7B 87 36.7 50.1 62.2
@@ 4.7 Inference Time Analysis
To evaluate the inference efficiency of the proposed ELMM approach, we compare its end-to-end inference time with that of state-of-the-art uni-modal methods based on large language models, including MKGL Guo et al. (2024a) and K-ON Guo et al. (2025b), using the FB15k-237-IMG dataset. As shown in Table 5, ELMM achieves lower inference time despite the need to process multimodal information, primarily due to its attention pruning strategy. This result highlights ELMM’s advantage in inference efficiency compared to uni-modal baselines. Moreover, by effectively modeling and integrating multimodal information, ELMM also achieves superior performance across standard evaluation metrics.
@@ 5 Conclusion
We propose Efficient Lightweight Multimodal Large Language Models (ELMM), a novel framework that extends Multimodal Large Language Models (MLLMs) to the task of Multimodal Knowledge Graph Completion (MKGC). ELMM addresses two key challenges: (1) semantic noise and modality conflicts caused by excessive image tokens, and (2) the substantial computational overhead associated with processing such inputs using standard MLLMs. To mitigate these issues, we propose a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism, which adaptively compresses image tokens from both textual and visual views. We further apply attention pruning to remove redundant attention layers from the MLLMs backbone, improving inference efficiency. To offset any performance degradation from pruning, we add a linear projection with a carefully designed initialization scheme. Experiments on two benchmark datasets, FB15k-237-IMG and WN18-IMG, show that ELMM achieves state-of-the-art performance.

### v2

@@ 3.5 Parameter Sensitivity Analysis
Figure 4: Hyperparameter analysis results. Left: Relationship between Hits@1 and the number of attention pruning layers. Right: Relationship between Hits@1 and number of image tokens are retained after MVTC.
@@ 3.5.1 Number of Pruning Layers
3.5.2 Number of Image Tokens are Retained After MVTC
We systematically evaluate the impact of the number of retained image tokens on model performance, as shown in Figure 4(right). Overall, performance improves as more image tokens are preserved, indicating the positive contribution of visual information to multimodal knowledge graph completion. However, when the number of retained image tokens exceeds 128, performance degrades, suggesting that excessive visual tokens may introduce redundancy or noise that hampers effective modeling. Balancing predictive performance and computational efficiency, we retain 64 visual tokens in subsequent experiments.
@@ 3.6 Impact of Backbone MLLMs on MKGC Performance
To investigate whether ELMM can continuously benefit from the rapid evolution of MLLMs, we study the impact of different backbone models on completion performance. Specifically, we select LLaVA-1.5 7B llava1.5, LLaVA-1.5 13B llava1.5, and Qwen2.5-VL 7B qwen2.5vl as comparative models. The results are presented in Table 4.
@@
LLaVA-1.5 7B 105 37.4 50.2 63.7
LLaVA-1.5 13B 102 37.8 50.7 64.0
Qwen2.5-VL 7B 83 38.6 52.2 64.9
@@ 3.7 Inference Time Analysis
We conduct experiments on the FB15k-237-IMG dataset to compare the end-to-end reasoning latency of ELMM with that of state-of-the-art unimodal large language model–based methods, including MKGL and K-ON. Specifically, we randomly sample 1,000 instances from the training set for inference, and report the average latency as the inference time. As shown in Table 5, despite the additional cost of processing multimodal inputs, ELMM exhibits lower inference latency than unimodal baselines. This efficiency gain is primarily attributed to the proposed attention pruning strategy, which effectively reduces redundant computation during reasoning. These results demonstrate that ELMM not only outperforms unimodal methods in reasoning efficiency, but also achieves strong performance on standard evaluation metrics by effectively modeling and integrating multimodal information.
@@ 4 Conclusion
We propose Efficient Lightweight Multimodal Large Language Models (ELMM), a novel framework that extends MLLMs to the task of MKGC. ELMM addresses two key challenges: (1) semantic noise and modality conflicts caused by excessive image tokens, and (2) the substantial computational overhead associated with processing such inputs using standard MLLMs. To mitigate these issues, we propose a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism, which adaptively compresses image tokens from both textual and visual views. We further apply attention pruning to remove redundant attention layers from the MLLMs backbone, improving inference efficiency. To offset any performance degradation from pruning, we add a linear projection with a carefully designed initialization scheme. Experiments on four benchmark datasets, show that ELMM achieves state-of-the-art performance.
Limitations
Although ELMM demonstrates promising effectiveness, it relies on powerful pretrained multimodal backbone models, such as LLaVA and Qwen2.5-VL. While our framework is largely model-agnostic, its performance and efficiency gains are still constrained by the representational capacity of the underlying multimodal large language models (MLLMs).
Appendix A License for Scientific Artifacts
The Training dataset FB15k-237-IMG FB15K-237 is licensed under Creative Commons Attribution 4.0 License [^1] https://choosealicense.com/licenses/cc-by-4.0/. The Training dataset WN18-IMG WN18 is licensed under WordNet Release 3.0 License [^2] https://wordnetcode.princeton.edu/3.0/README. The Training dataset DB15K DB15K is licensed under CC BY-SA 3.0 License [^3] https://creativecommons.org/licenses/by-sa/3.0/deed.en. The Training dataset MKG-W MMRNS is licensed under CC BY-SA 3.0 License [^4] https://creativecommons.org/licenses/by-sa/3.0/deed.en. The LLaVA-1.5 model llava1.5 and Qwen2.5-VL model qwen2.5vl are licensed under Apache License 2.0555https://choosealicense.com/licenses/apache-2.0/. All usages of scientific artifacts in this paper obey the corresponding licenses.
Appendix B Related Works
B.1 Multimodal Knowledge Graph Completion
Multimodal knowledge graph completion (MKGC) seeks to enhance the reasoning ability of knowledge graphs by jointly leveraging structural triples and multimodal entity information such as images and text. Early approaches TransAE; RSME; IKRL typically used simple feature concatenation, but recent research has introduced more advanced techniques for fine-grained and adaptive multimodal fusion. Representative models include MKGformer MKGformer, which employs hybrid transformers with multi-level fusion; SGMPT SGMPT introduces structural information based on MKGformer; LAFA LAFA, which adopts link-aware fusion and neighbor aggregation; MyGO MyGO, which tokenizes multimodal entity information for fine-grained representation; and NativE zhang2024native, which addresses modality imbalance in real-world scenarios.
With the rapid advancement of multimodal large language models (MLLMs), these models have demonstrated substantial potential in task related to Multimodal Knowledge Graphs (MKGs). However, most studies have yet to fully explore and exploit the capabilities of MLLMs. To address this gap, we propose ELMM, a novel method designed to harness the strengths of MLLMs.
Datasets MLLM LoRA r LoRA dropout LoRA Target
FB15k-237-IMG LLaVA-1.5 7B 32 0.05 query, value
WN18-IMG LLaVA-1.5 7B 32 0.05 query, value
DB15K LLaVA-1.5 7B 32 0.06 query, value
MKG-W LLaVA-1.5 7B 32 0.06 query, value
Table 6: LoRA settings in the main experiments.
Datasets Layer of Pruning
FB15k-237-IMG 13,24,28,20,22,29,12,30,9,19,23,21,18,31,5,26
WN18-IMG 27,17,9,20,13,23,30,6,31,14,25,22,21,26,28,19
DB15K 15,27,10,19,17,13,20,29,21,22,31,30,28,23,25,14
MKG-W 19,8,17,16,27,30,23,11,22,29,20,24,21,26,18,31
Table 7: Attention pruning settings in the main experiments.
B.2 Multimodal Large Language Models
With the rapid development of Multimodal Large Language Models (MLLMs), such as GPT-4 achiam2023gpt, LLaVA-1.5 llava1.5, Qwen2.5-VL qwen2.5vl, and Emu2 Emu2, vision and language understanding have been deeply integrated within a unified semantic space. These models typically adopt a paradigm where visual encoders vit segment images into multiple regional patches, which are then encoded into dense image token sequences. These image tokens are projected to align with the dimensionality of text tokens and concatenated with textual inputs for joint processing by transformer backbones, enabling strong performance in visual question answering, image captioning, and general knowledge reasoning tasks.
In recent years, the redundancy of visual tokens has been widely recognized as a major source of increased computational and memory costs in multimodal models, prompting growing interest in addressing the inefficiency caused by excessive image tokens Deco; IPCV; wen2025stop; yang2025efficientvlatrainingfreeaccelerationcompression. For instance, Qwen2.5-VL qwen2.5vl incorporates dynamic resolution adjustment, DeCo Deco performs parameter-free adaptive pooling compression at the patch level, and IPCV IPCV mitigates visual token redundancy through shallow-layer token pruning and neighbor-guided reconstruction for recovering pruned tokens.
However, existing methods compress visual tokens solely based on visual view, which is not well suited for the MKGC task. As demonstrated in Figure 1 and further validated by our ablation studies, visual tokens that are semantically aligned with the textual modality play a more critical role in MKGC, a factor largely overlooked by prior approaches. To address this limitation, we propose MVTC, which leverages textual information to identify and preserve the most relevant visual tokens while compressing those that are less informative. This design facilitates more effective multimodal alignment and fusion in subsequent reasoning stages. Furthermore, to enable the application of MLLMs to MKGC, ELMM proposes several additional innovations beyond MVTC, including an attention pruning strategy and a multimodal knowledge reasoning completion layer. Together, these components allow ELMM to achieve state-of-the-art performance on MKGC benchmarks.
Appendix C Further Experiment Details
C.1 Training Hyperparameters.
We adopt LLaVA-1.5 7B llava1.5 as the base Multimodal Large Language Model (MLLM), incorporating Low-Rank Adaptation (LoRA) hu2022lora techniques into both the query and value layers to enable parameter-efficient fine-tuning. For LoRA related parameter settings, please refer to Table 6. All experiments are conducted using eight NVIDIA A100 GPUs, and both training and inference are implemented with the PyTorch framework. The parameter $K$ in the attention pruning strategy is always set to 16, indicating that 16 attention modules in the MLLM are pruned and replaced with linear projections. To determine which layers to prune, we randomly sample 1,000 training instances and compute the cosine similarity between the attention maps of adjacent layers for each dataset. Layers exhibiting the highest similarity—indicating strong redundancy—are selected for pruning. The specific pruned layers for each dataset are reported in Table 7. The key parameter $H$ in MVTC represents the number of attention headers, and $H$ is set to 32 in all datasets, which represents 32 image tokens each for the textual and visual views, resulting in a total of 64 image tokens. For training parameter settings, please refer to Table 8. We report the version numbers of used packages in Table 9.
Datasets MLLM Batch Size Optimizer Epoch Learning Rate
FB15k-237-IMG LLaVA-1.5 7B 32 Adam 5 3e-4
WN18-IMG LLaVA-1.5 7B 16 Adam 2 2e-4
DB15K LLaVA-1.5 7B 64 Adam 2 1e-4
MKG-W LLaVA-1.5 7B 64 Adam 2 1e-4
Table 8: Training settings in the main experiments.
C.2 Prompt used in the experiment.
Across all datasets, we employ a unified training and evaluation prompt. Specifically, for instances where the task is to predict the tail entity given a head entity and a relation, we use the following prompt formulation:
    "Next you will be given the Head entity,
    which has image representations:
    <image>.
    Suppose that you are an excellent linguist
    studying a three-word language.
    Given the following dictionary:
[…]
    Please complete the last word (?)
    of the sentence: {h}{r}?"
Similarly, for instances where the task is to predict the head entity given a relation and a tail entity, we use the following prompt formulation:
    "Next you will be given the Head entity,
    which has image representations:
    <image>.
    Suppose that you are an excellent linguist
    studying a three-word language.
    Given the following dictionary:
[…]
    Please complete the last word (?)
    of the sentence: {t}{inv_r}?"
Package Version Package Version
PyTorch 2.0.0 transformers 4.48.0
deepspeed 0.10.0 tokenizers 0.20.1
scatter 2.1.2 sparse 0.6.18
datasets 2.14.3
Table 9: Versions of used packages.
Model Name DB15K MKG-W
MRR Hits@1 Hits@3 Hits@10 MRR Hits@1 Hits@3 Hits@10
Baseline Models
IKRL(IJCAI 2017) 26.8 14.1 34.9 49.1 32.4 26.1 34.8 44.1
TBKGC(NAACL 2018) 28.4 15.6 37.0 49.9 31.5 25.3 34.0 43.2
TransAE(IJCNN 2019) 28.1 21.3 31.2 41.2 30.0 21.2 34.9 44.7
MMKRL(APIN 2025) 26.8 13.9 35.1 49.4 30.1 22.2 34.1 44.7
RSME(MM 2021) 29.8 24.2 32.1 40.3 29.2 23.4 32.0 40.4
VBKGC(KDD 2022) 30.6 19.8 37.2 49.4 30.6 24.9 33.0 40.9
OTKGE(NeurIPS 2022) 23.9 18.5 25.9 34.2 34.4 28.9 36.3 44.9
IMF(WWW 2023) 32.3 24.2 36.0 48.2 34.5 28.8 36.6 45.4
QEB(MM 2023) 28.2 14.8 36.7 51.6 32.4 25.5 35.1 45.3
VISTA(EMNLP 2023) 30.4 22.5 33.6 45.9 32.9 26.1 35.4 45.6
AdaMF(COLING 2024) 32.5 21.3 39.7 51.7 34.3 27.2 37.9 47.2
MyGO(AAAI 2025) 37.7 30.1 41.3 52.2 36.1 29.8 38.5 47.8
MCKGC(AAAI 2025) 39.8 31.9 43.8 54.7 36.9 31.3 38.9 47.4
K-ON (AAAI 2025) 38.1 30.1 42.8 53.6 36.6 30.1 38.7 48.3
ELMM(Ours) 41.2 34.1 45.5 56.7 38.4 33.5 41.7 51.5
Table 10: Models performance on the DB15K and MKG-W datasets. Bold and underline text indicate the best and second-best performance, respectively.
Appendix D More Experiment Results.
D.1 Main Results on DB15K and MKG-W.
D.1.1 Comparison Methods and Evaluation Protocol.
Since many methods evaluated on FB15k-237-IMG and WN18-IMG are not tested on DB15K and MKG-W, and vice versa, we re-select the comparison methods to enable a fair and meaningful evaluation on the DB15K and MKG-W datasets. The selected baselines primarily include IKRL IKRL, TBKGC TBKGC, TransAE TransAE, MMKRL MMKRL, RSME RSME, VBKGC vbkgc, OTKGE OTKGE, IMF IMF, QEB QEB, VISTA VISTA, AdaMF AdaMF-MAT,MCKGC MCKGC, MyGO MyGO,and K-ON kno.
We adopt standard evaluation metrics, including Hits@k (Hits@1, Hits@3, and Hits@10) to measure ranking accuracy at different cutoffs, and Mean Reciprocal Rank (MRR) to evaluate the overall ranking performance of each model.
D.1.2 Main results.
We conduct a comprehensive comparison between the proposed ELMM and 14 state-of-the-art methods on the DB15K and MKG-W datasets. As reported in Table 10, ELMM consistently outperforms all competing approaches across all evaluation metrics, which is in line with the results observed on the FB15k-237-IMG and WN18-IMG datasets. These results demonstrate the robustness and general effectiveness of ELMM for multimodal knowledge graph completion, and further highlight the potential of multimodal large language models (MLLMs) in advancing multimodal knowledge graph completion.
D.2 Robustness Experimental Test.
We conducted a systematic evaluation of the robustness of ELMM under modality-missing scenarios. Specifically, to simulate modality incompleteness in real-world applications, we randomly removed either textual descriptions or visual information from a subset of training samples, while keeping the evaluation procedures unchanged. This setting is designed to assess the ability of ELMM to operate effectively under incomplete multimodal inputs.We considered modality-missing rates of 20%, 40%, 60%, and 80%. For example, a 20% missing rate indicates that 10% of the training samples lack textual information and an additional 10% lack visual information; higher missing rates follow the same proportion.
All experiments were conducted on the FB15k-237-IMG dataset, with Hits@1 adopted as the evaluation metric. The experimental results are illustrated in Figure 5. An interesting phenomenon can be observed: when the modality missing rate is set to 20%, the model achieves a noticeable performance improvement compared to the fully observed setting. We hypothesize that a moderate degree of modality missing increases the training difficulty, thereby exerting a regularization effect that encourages the model to learn more robust representations.
As the modality missing rate further increases, the overall Hits@1 performance exhibits a downward trend. Nevertheless, even under relatively high missing rates, the performance degradation remains limited. This observation indicates that ELMM demonstrates strong robustness in modality-missing scenarios. We attribute this robustness to the ability of multimodal large language models to leverage their internal knowledge and contextual reasoning capabilities to partially compensate for missing modality information, thereby mitigating the adverse effects caused by modality incompleteness.
Figure 5: Performance under varying modality-missing rates on FB15k-237-IMG. The y-axis corresponds to Hits@1, and the x-axis represents the proportion of missing modality information.